OpenStudy (anonymous):

evaluate this indefinite integral:

OpenStudy (anonymous):

\[\int\limits_{-1}^{5} \left| x-4x^2 \right|\]

OpenStudy (anonymous):

best figure out what your piecewise function really is before you start

OpenStudy (anonymous):

I tried splitting it up into two equations but got confused

OpenStudy (anonymous):

first you have to solve \(x-4x^2>0\)

OpenStudy (anonymous):

whoops, definite integral, my post said indefinite

OpenStudy (anonymous):

Yeah, so would I just say x>4x^2?

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

\(x-4x^2\) is a parabola that opens down it is 0 at \(x=0\) and at \(x=\frac{1}{4}\) so it is positive between the zeros, namely on \((0,\frac{1}{4})\) and negative everywhere else

OpenStudy (anonymous):

here is the picture with the solution http://www.wolframalpha.com/input/?i=x-4x^2%3E0

OpenStudy (anonymous):

so I should probably do the alternate one so that it's positive right?

OpenStudy (anonymous):

thefore you have to break the integral up in to 3 (!) pieces \[\int_{-1}^0(4x^2-x)dx\] \[\int_0^{\frac{1}{4}}(x-4x^2)dx\] \[\int_{\frac{1}{4}}^5(4x^2-x)dx\]

OpenStudy (anonymous):

alright, I think I can do it from there, thanks!

OpenStudy (anonymous):

yw